2: An aerosol can contains gases under a pressure
of 4.6 atm at 23 ◦C. If the can is left on a
hot sandy beach, the pressure of the gases
increases to 4.78 atm. What is the Celsius
temperature on the beach?
Answer in units of ◦C
4: The volume of a gas is 22.6 mL at 46◦C and
0.82 atm. What will the volume be at 15◦C
and 0.996 atm?
Answer in units of mL. 5: The volume of a gas at 25◦C and 0.29 atm
is 83 mL. What volume will the same gas
sample occupy at standard conditions?
Answer in units of mL.
6: A sample of an ideal gas occupies 215 mL at
31.2
◦C and 714 torr.
What is its volume at STP?
1
Expert's answer
2020-04-20T15:09:15-0400

2)

"\dfrac{p_1}{T_1}=\dfrac{p_2}{T_2}\\\dfrac{4.6}{296}=\dfrac{4.78}{T_2}"

"T_2=" 307.6

in Celsius it is (307.6-273)=33.6oC